You push a 55 N box 21.5 meters up a 15.2 degree incline. The coefficient of friction with the floor is 0.212 against which you apply a constant 27.7 N of force. How much Work have you done, in Joules?
Answer #1

Work done is the multiple of force and displacement in the direction of force applied.

So given,

Displacement (d) = 22.5 m

Force (F) = 27.7 N

Work done (W ) = F × d = 22.5×27.7 Jules

W = 623.25 J

thank you please rate the solution.